# Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

### v6

@@ 6.1 Exact topological Euler link
Combining with Proposition 6.2 yields the identity
$\boxed{\,T=V-3+S_{i}\,}.$
6.2 Heuristic incidence formulas for external structure
The next statement is not a general topological law; it is a combinatorial incidence calculation that assumes:
@@ 6.2 Heuristic incidence formulas for external structure
    all internal adjacencies are along triangular faces,
@@
    $S$ counts boundary coplanarity merges,
@@
Assume a genus-zero decomposition into $T$ tetrahedra without interior Steiner vertices, where internal adjacencies occur only along triangular faces; let $N_{i}$ be the number of internal gluing triangles, $S$ the number of coplanarity merges on the boundary, and let $S_{i}$ denote the number of interior edges.
Then the boundary satisfies
@@
For a fully triangulated boundary, $F_{\triangle}=4T-2N_{i}$ and $E_{\triangle}=\frac{3}{2}F_{\triangle}=6T-3N_{i}$; $S$ coplanarity merges reduce both by $S$, giving the displayed $E$ and $F$. The $V$-formula is algebraically equivalent to the exact identities $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (Propositions 6.26.3), hence consistent. ∎
Consistency check.
Applying these expressions to two distinct polyhedra with $V=9$ and $S=5$:
Domed octagon ($T=6$, $N_{i}=5$, $S=5$, $S_{i}=0$):
$E=6\cdot 6-3\cdot 5-5=16,\quad F=4\cdot 6-2\cdot 5-5=9,\quad V=4\cdot 6-3\cdot 5+0=9.$
Cube with pyramid ($T=7$, $N_{i}=7$, $S=5$, $S_{i}=1$):
$E=6\cdot 7-3\cdot 7-5=16,\quad F=4\cdot 7-2\cdot 7-5=9,\quad V=4\cdot 7-3\cdot 7+2=9.$
In both cases: $E=16$, $F=9$, $V=9$, confirming consistency.
6.3 Heuristic extension of Euler’s characteristic
Substituting the above heuristic expressions into VE+FV-E+F gives:
$V-E+F=(4T-3N_{i}+2S_{i})-(6T-3N_{i}-S)+(4T-2N_{i}-S)=2(T-N_{i}+S_{i}).$
Heuristic 6.1 (Extended Euler-type identity).
@@ 6.3 Heuristic extension of Euler’s characteristic
Note: This is an incidence calculation that packages the boundary formulas above; it is not asserted as a general topological law. In normal form, $T-N_{i}+S_{i}=1$ (Proposition 6.3), hence VE+F=2V-E+F=2 as expected.
Reduction to the classical value in a specific family.
For the sequence $T_{k}=V-3+k$, $N_{i,k}=V-4+2k$, $S_{i,k}=k$ ($0\leq k\leq V-5$), one finds $T_{k}-N_{i,k}+S_{i,k}=1$, hence VE+F=2V-E+F=2.
@@ 7 A Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility
Then for any genus-0 polyhedron in normal form,
@@ 7.0.1 Derivation (for completeness)
Step 4: Triangulation Segment Count (S):
Count non-intersecting segments used in triangulation of faces:
$S_{\operatorname{total}}=\sum_{k=3}^{n}(k-3)P_{k}$
@@
Step 6: Edge Count Estimate (E):
$E=3(V-2)-S_{\operatorname{total}}$
Equivalently, this identity can be rearranged to express a relationship between the total edge incidences and the number of triangulated faces: $2E+S_{\operatorname{total}}=3F$
This confirms that the triangulated surface has exactly $3F$ triangle-edge incidences, spread across edges and triangulation segments.
@@
$S_{\min}=0$
$S_{\max}=\left\{\begin{array}[]{ll}\frac{3}{2}V-\tfrac{13}{2}&\text{if }V\text{is odd}\\ &\\ \frac{3}{2}V-6&\text{if }V\text{is even}\end{array}\right$.
$S_{\operatorname{difference}}=S_{\max}-S_{\operatorname{total}}$
If $S_{\operatorname{difference}}<0$, the set of polygons is “too flat” to enclose a 3D volume.
@@
The combinatorial method introduced in this paper has been applied to a range of polygonal configurations, demonstrating its capacity to correctly validate known polyhedra, reject over-flattened ones, and expose structurally deceptive false positives. These worked examples include the cube, the square pyramid, and the regular dodecahedron, each of which passes all symbolic checks and matches known geometry. For example, the decomposition of the cube yields five tetrahedra which aligns with a known meshing scheme producing four right tetrahedra at its corners and a central regular tetrahedron [7]. In contrast, a configuration of 36 hexagons fails the flatness threshold, while a less intuitive case, two hexagons and four triangles, passes all algebraic constraints but appears unrealizable in $\mathbb{R}^{3}$, highlighting the symbolic method’s limits. Intriguingly, substituting two triangles with one square in that borderline case causes the structure to fail the flatness test entirely, revealing a sharp combinatorial threshold. Such substitutions may serve as diagnostic probes for detecting “brittleness” in enclosure feasibility. These examples reinforce the method’s utility as both a filtering tool for polyhedral candidates and a lens through which to explore the fine boundary between combinatorial sufficiency and geometric necessity.
@@ 8.1 Face Types and Vertex Types
where $d_{i}$ is the degree of the $i$-th face and $v_{j}$ is the valency of the $j$-th vertex. This face–vertex symmetry is emphasized in Grünbaum’s treatment of polyhedral duality [6], where vertex valencies and face degrees are shown to reflect a deeper structural reciprocity across convex forms. It reflects a basic balance: each edge is shared by two faces and connects two vertices, hence contributes twice to the totals.
@@ 8.3 Face-Type Combinatorial Constraints
The following determines valid face-type combinations in a polyhedron defined by:
  • Total number of vertices $V$,
  • External triangulation segment count $S$,
  • Total number of faces $F$,
  • Total number of edges $E$.
Face Variables and Degree Constraints.
Let the variables represent the number of faces of each type:
  • $a$: triangles (3 sides),
  • $b$: quadrilaterals (4 sides),
  • $c$: pentagons (5 sides),
  • etc., up to degree V1V-1.
Face Type Limit Rule. The face with the largest number of edges allowed in a polyhedron with $V$ vertices must have at most V1V-1 edges, since a face cannot be incident with more than V1V-1 unique vertices in a closed 3D structure.
@@
$a+b+c+\ldots$ $=F$ (face count)
$3a+4b+5c+\ldots$ $=2E$ (edge incidence)
$0a+1b+2c+\ldots$ $=S$ (triangulation segments).
@@
  • $a=4$, $b=2$, $c=0$  (4 triangles, 2 quadrilaterals),
  • $a=5$, $b=0$, $c=1$  (5 triangles, 1 pentagon).
@@ 8.4 Vertex-Type Combinatorial Constraints
We now narrow down the valid vertex-type combinations in a polyhedron defined by:
  • $V$: total number of vertices,
  • $E$: total number of edges,
  • $S$: number of triangulation segments,
  • $T$: total number of tetrahedra.
Vertex Variables.
Let the variables represent the number of vertices of each degree type:
  • $a^{\prime}$: vertices of degree 3,
  • $b^{\prime}$: vertices of degree 4,
  • $c^{\prime}$: vertices of degree 5,
  • etc., up to degree V1V-1.
@@
$a^{\prime}+b^{\prime}+c^{\prime}+\ldots$ $=V$ (vertex count)
$3a^{\prime}+4b^{\prime}+5c^{\prime}+\ldots$ $=2E$ (edge incidence)
$a^{\prime}+2b^{\prime}+3c^{\prime}+\ldots$ $=2E-2V$ (since $\sum_{j}(\operatorname{deg}v_{j}-2)=2E-2V$).
@@
  • $a^{\prime}=4$, $b^{\prime}=2$, $c^{\prime}=0$,
  • $a^{\prime}=5$, $b^{\prime}=0$, $c^{\prime}=1$.
@@ 8.5 Limitations and Realizability Filters
This illustrates a key limitation: satisfying combinatorial constraints does not guarantee embeddability in three-dimensional space. As Grünbaum notes, only graphs that are planar and 3-connected can be realized as convex 3D polyhedra, a result due to Steinitz [6].
@@ 9 On All Likely Configurations of a Polyhedron and The Limits of Combinatorial Realizability
I now define a function that estimates the number of valid combinations between these two sets for a given vertex count $V$. This count reflects how many internally and externally compatible sets of structures exist, without specifying the actual types of those structures (Face-types and Vertex-types).
9.1 Structure Pair Combinations
@@
$\text{Combinations}(V)=\left\{\begin{array}[]{ll}\frac{1}{2}V^{2}-3V+5&\text{if }V\text{is even}\\ \frac{1}{2}V^{2}-3V+\tfrac{9}{2}&\text{if }V\text{is odd}\end{array}\right$.
9.2 Example Table (V = 4 to 20)
$V$ (Vertices) Estimate Total External and Internal Configuration Sets
4 1
5 2
6 5
7 8
8 13
9 18
10 25
11 32
12 41
13 50
14 61
15 72
16 85
17 98
18 113
19 128
20 145
9.2.1 Graphical Representation
Figure 4: Combinatorial sum of internal and external structure groupings as a function of vertex count.
@@ 9.2.1 Graphical Representation
When attempting to uncover the actual face-type and vertex-type configurations corresponding to a specific polyhedron defined by vertex count $V$, triangulation segment count $S$, and tetrahedron count $T$, the procedures described in Section 7 (Constraints on Face Type Integer Combinations and Maximum Face Degree Rule) and Section 8 (Constraints on Vertex Type Integer Combinations and Maximum Edges per Vertex Degree Rule) should be applied. These methods allow one to solve for specific polygonal face arrangements and converge to vertex degree structures that are compatible with a given global combinatorial configuration.
@@ 9.3 Estimating the Upper Bound from Flatness and Face Combinations
While previous sections emphasize compatibility between internal and external combinatorics, it is also useful to estimate the maximum number of possible polyhedral configurations based solely on the total surface angle budget. In contrast to a full combinatorial analysis of vertex and edge constraints, this method estimates the number of ways one can arrange polygonal faces on a polyhedral surface using only the flatness range of the surface. It does not take into account local realizability constraints, such as vertex incidence or internal fillability by tetrahedra. As it doesn’t account for vertex types, it serves as a partial initial upper-bound estimate of surface complexity.
@@ 9.4 Method Outline
Define the triangulation segment budget $S$, where $S=0$ corresponds to a surface composed entirely of triangles (i.e., zero extra flatness beyond minimal curvature), and larger values of $S$ correspond to the inclusion of increasingly flat faces (e.g., quadrilaterals, pentagons). The minimum is always $S_{\min}=0$, corresponding to the fully triangulated case. The maximum possible number of triangulation-equivalent segments $S_{\max}$ depends on the number of vertices $V$ and follows the previously mentioned empirically derived linear functions:
$S_{\max}=\left\{\begin{array}[]{ll}\frac{3}{2}V-\tfrac{13}{2}&\text{if }V\text{is odd}\\ \frac{3}{2}V-6&\text{if }V\text{is even}\end{array}\right$.
@@
Thus, every value of $S\in[0,S_{\max}]$ corresponds to a number of integer partitions under these constraints. The total number of valid polygonal configurations at a given $V$ is then equal to the sum of these partition counts from $S_{\min}$ to $S_{\max}$. If no restriction is placed on the number of polygon sides, then this count becomes equivalent to the unrestricted partition function $p(S)$, a classical object in number theory that enumerates the number of ways to write $S$ as a sum of positive integers (ignoring order) [2]. The partition function begins:
@@
$\sum_{S=0}^{S_{\max}}p(S)$
For instance, for $V=5$, I compute $S_{\max}=1$, and the number of valid configurations is:
@@
Vertices $S_{\min}$ $S_{\max}$ Face Type Combinations (Upper Bound)
@@ 9.5 Graphical Representation
If a maximum polygon side count $k_{\max}$ is imposed (such as limiting the use of hexagons or excluding higher-order polygons) the enumeration becomes a case of restricted integer partitions, where only parts of size $\leq k_{\max}-3$ are allowed. This constraint reflects that each non-triangular face contributes a specific number of excess triangulation segments. Restricted partitions offer a more realistic estimate by excluding configurations that rely on geometrically unlikely face types. For example, limiting the count of hexagons in some configurations prevents overcounting face combinations that cannot physically tile a polyhedral surface. By summing the number of valid restricted partitions for each value of $S=0$ to $S_{\max}$, I obtain a tighter upper bound on the number of feasible face-type combinations, filtering out many false positives that are combinatorially valid but geometrically unconstructible.

### v7

@@ 5.8.1 Exact Euler link and normal-form specialization
Moreover, after triangulating all boundary faces, triangle incidence gives (Proposition 5.2)
$\boxed{\,N_{i}=2T-V+2\,}.$
Together these identities link the interior parameters $(T,N_{i},S_{i})$ to the boundary vertex count $V$ in a purely symbolic way.
5.8.2 Heuristic incidence formulas for external structure
The next statement is not a general topological law; it is an incidence calculation that assumes:
@@ 5.8.2 Heuristic incidence formulas for external structure
    all internal adjacencies occur along triangular faces,
@@
    $S=\sum_{f}(\deg f-3)$ is the face-degree excess (equivalently the number of boundary triangulation diagonals),
@@
Assume a genus–0 decomposition into $T$ tetrahedra without interior Steiner vertices, where internal adjacencies occur only along triangular faces; let $N_{i}$ be the number of internal gluing triangles, $S$ the number of coplanarity merges on the boundary, and let $S_{i}$ denote the number of interior edges. Then the boundary satisfies
@@
For a fully triangulated boundary, $F_{\triangle}=4T-2N_{i}$ and $E_{\triangle}=\tfrac{3}{2}F_{\triangle}=6T-3N_{i}$ by triangle incidence. Coplanarity merges reduce both $E_{\triangle}$ and $F_{\triangle}$ by $S$, yielding the displayed formulas for $E$ and $F$. The VV–formula is algebraically equivalent to the exact identities $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (Propositions 5.25.3), hence consistent. ∎
Sanity check.
In normal form, $T-N_{i}+S_{i}=1$ (Proposition 5.3), so the above formulas automatically satisfy VE+F=2V-E+F=2 after substitution.
5.8.3 Heuristic extension of Euler’s characteristic
Substituting the heuristic expressions of Proposition 5.4 into VE+FV-E+F gives
$V-E+F=2\,(T-N_{i}+S_{i}).$
Heuristic 5.2 (Extended Euler-type identity).
@@ 5.8.3 Heuristic extension of Euler’s characteristic
Note: This is an incidence packaging of the boundary formulas above; it is not asserted as a general topological law. In normal form, $T-N_{i}+S_{i}=1$ (Proposition 5.3), hence VE+F=2V-E+F=2 as expected.
Appendix F compiles all identities and bounds derived in Sections 4 and 5.
@@ 6 A Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility
Then for any finite simple graph with a 2–cell embedding in $S^{2}$ (genus 0),
@@
Step 4: Face-degree excess (flatness) $S$:
$S:=\sum_{k\geq 3}(k-3)\,P_{k}.$
@@
Step 6: Edge count (E):
E=3V6S.E=3V-6-S.
Equivalently, after triangulation one has $F_{\triangle}=F+S$ and $E_{\partial}=E+S$, and the triangulated incidence identity
$2E_{\partial}=3F_{\triangle}\quad\text{ i.e. }\quad 2(E+S)=3(F+S)$
holds.
@@
$S_{\mathrm{lb}}=0$
$S_{\mathrm{ub}}=\left\{\begin{array}[]{ll}\frac{3}{2}V-\tfrac{13}{2}&\text{if }V\text{is odd}\\ &\\ \frac{3}{2}V-6&\text{if }V\text{is even}\end{array}\right$.
$S_{\mathrm{diff}}=S_{\mathrm{ub}}-S$
If $S>S_{\mathrm{ub}}(V)$, the face-degree excess exceeds the parity/degree upper bound implied by $\delta(G)\geq 3$, so no such genus-zero polyhedral graph can exist. Loosely, one may say the proposed surface is “too flat (combinatorially) to enclose”.
@@
The combinatorial method introduced in this paper has been applied to a range of polygonal configurations, demonstrating its capacity to correctly validate known polyhedra, reject over-flattened ones, and expose structurally deceptive false positives. These worked examples include the cube, the square pyramid, and the regular dodecahedron, each of which passes all symbolic checks and matches known geometry. For example, the decomposition of the cube yields five tetrahedra which aligns with a known meshing scheme producing four right tetrahedra at its corners and a central regular tetrahedron [naff2015]. In contrast, a configuration of 36 hexagons fails the flatness threshold, while a less intuitive case, two hexagons and four triangles, passes all algebraic constraints but appears unrealizable in $\mathbb{R}^{3}$, highlighting the symbolic method’s limits. Intriguingly, substituting two triangles with one square in that borderline case causes the structure to fail the flatness test entirely, revealing a sharp combinatorial threshold. Such substitutions may serve as diagnostic probes for detecting “brittleness” in enclosure feasibility. These examples reinforce the method’s utility as both a filtering tool for polyhedral candidates and a lens through which to explore the fine boundary between combinatorial sufficiency and geometric necessity.
@@ 7.1 Face Types and Vertex Types
where $d_{i}$ is the degree of the $i$-th face and $v_{j}$ is the valency of the $j$-th vertex. This face–vertex symmetry is emphasized in Grünbaum’s treatment of polyhedral duality [gruenbaum2003], where vertex valencies and face degrees are shown to reflect a deeper structural reciprocity across convex forms. It reflects a basic balance: each edge is shared by two faces and connects two vertices, hence contributes twice to the totals.
@@ 7.3 Face-Type Combinatorial Constraints
We now narrow down the valid face-type combinations in a polyhedron vertex count $V$, edge count $E$, face count $F$, and external flatness $S$.
Face-Type Limit Rule.
In a polyhedral graph with $V$ vertices, any face can have at most V1V-1 edges, since its boundary is a simple cycle and therefore cannot visit all $V$ vertices without repetition.
@@
$F_{3}+F_{4}+F_{5}+\cdots$ $=F$ $\text{(total number of faces)},$
$3F_{3}+4F_{4}+5F_{5}+\cdots$ $=2E$ $\text{(edge--face incidence)},$
$(3{-}3)F_{3}+(4{-}3)F_{4}+(5{-}3)F_{5}+\cdots$ $=S$ $\text{(external triangulation segments)}.$
@@
  • $F_{3}=4$, $F_{4}=2$, $F_{5}=0$  (4 triangles, 2 quadrilaterals),
  • $F_{3}=5$, $F_{4}=0$, $F_{5}=1$  (5 triangles, 1 pentagon).
@@ 7.4 Vertex-Type Combinatorial Constraints
We now narrow down the valid vertex-type combinations in a polyhedron with total vertex count $V$ and edge count $E$.
@@
$V_{3}+V_{4}+V_{5}+\ldots$ $=V$ (vertex count)
$3V_{3}+4V_{4}+5V_{5}+\ldots$ $=2E$ (edge incidence)
$V_{3}+2V_{4}+3V_{5}+\ldots$ $=2E-2V$ (since $\sum_{j}(\operatorname{deg}v_{j}-2)=2E-2V$).
@@
  • $V_{3}=4$, $V_{4}=2$, $V_{5}=0$,
  • $V_{3}=5$, $V_{4}=0$, $V_{5}=1$.
@@ 7.5 Limitations and Realizability Filters
This illustrates a key limitation: satisfying combinatorial constraints does not guarantee embeddability in three-dimensional space. As Grünbaum notes, only graphs that are planar and 3-connected can be realized as convex 3D polyhedra, a result due to Steinitz [gruenbaum2003].
@@ 8 On All Likely Configurations of a Polyhedron and The Limits of Combinatorial Realizability
I now define a function that estimates a lower bound on the number of valid combinations between these two sets for a given vertex count $V$. We denote this conservative estimate by $\mathrm{Combinations}_{\min}(V)$. This count reflects how many internally and externally compatible sets of structures exist, without specifying the actual types of those structures (face-types and vertex-types). Refining this lower bound by characterizing the persistence of internal multiplicity as a function of $S$ remains an open combinatorial problem.
8.1 Conservative Estimate for Structure Pair Combinations
@@
$\mathrm{Combinations}_{\mathrm{lb}}(V)=\begin{cases}\frac{1}{2}V^{2}-3V+5,&\text{if $V$ is even},\\[4.0pt] \frac{1}{2}V^{2}-3V+\tfrac{9}{2},&\text{if $V$ is odd}.\end{cases}$
Derivation.
Fix a vertex count $V\geq 4$. The external flatness parameter $S$ ranges over
$S=0,1,\dots,S_{\mathrm{ub}}(V),\qquad S_{\mathrm{ub}}(V)=3V-6-\left\lceil\frac{3V}{2}\right\rceil=\begin{cases}\frac{3V}{2}-6,&V\ \text{even},\\[2.0pt] \frac{3V}{2}-\frac{13}{2},&V\ \text{odd}.\end{cases}$
For each fixed $S$, the internal tetrahedral count is constrained by the internal ladder bounds
$T_{\min}(V)=V-3\qquad\text{and}\qquad T_{\max}(V,S)=2V-8-S.$
Thus, at the level of global internal invariants, the number of admissible integer values of $T$ compatible with a given $S$ is at most
$T_{\max}(V,S)-T_{\min}(V)+1=\bigl(2V-8-S\bigr)-(V-3)+1=V-4-S.$
In particular, as $S$ increases, the admissible $T$-range shrinks linearly. At $S=0$ (fully triangulated boundary), there exist polyhedra that realize the full ladder of internal invariant values $(T,N_{i},S_{i})$, with the bipyramid–prism family attaining the maximal value $T_{\max}$.
To obtain a conservative macro-level pairing count, we require only that at least one internally compatible invariant triple exists for each admissible value of $S$, even after the $T$-interval collapses. Writing this as a capped count gives
$N_{\mathrm{int}}(V\mid S)=\max\{\,V-4-S,\;1\,\}=\max\{\,2V-8-S-(V-3)+1,\;1\,\}.$
Therefore the total number of compatible external–internal macro-pairs is
$\sum_{S=0}^{S_{\mathrm{ub}}}N_{\mathrm{int}}(V\mid S)=\sum_{S=0}^{V-5}(V-4-S)+\sum_{S=V-4}^{S_{\mathrm{ub}}}1.$
The first term is a triangular sum,
$\sum_{S=0}^{V-5}(V-4-S)=(V-4)+(V-5)+\cdots+1=\frac{(V-4)(V-3)}{2},$
and the second term contributes a linear tail of length $S_{\mathrm{ub}}-(V-4)+1$. Substituting the parity form of $S_{\mathrm{ub}}$ and simplifying yields
$\mathrm{Combinations}_{\mathrm{lb}}(V)=\begin{cases}\frac{1}{2}V^{2}-3V+5,&V\text{ even},\\[4.0pt] \frac{1}{2}V^{2}-3V+\tfrac{9}{2},&V\text{ odd}.\end{cases}$
Here the quadratic leading term arises from summing the linearly shrinking internal ladder across $S$, while any multiplicity of realizations for a given $(V,S,T)$ affects only lower-order corrections beyond this macro-invariant count.
It is worth emphasizing again that this estimate is intentionally conservative. As illustrated in Table 2 for $V=8$, multiple distinct internal configurations may persist even after the nominal $T$-interval has ceased to shrink linearly with $S$. Such persistence increases the true number of compatible structure pairs beyond the value predicted by $\text{Combinations}(V)$ and indicates that the present formula provides a lower bound rather than an exact count.
Remark (tail persistence and decomposition pathways).
Table 2 shows that the “tail” regime in $S$ need not collapse to a single internal invariant triple. For example, at $V=8$ the cube admits a $T=5$ decomposition obtainable by successive “shaving” moves, but also admits $T=6$ decompositions obtained by first partitioning the volume into two hexahedral blocks and then tetrahedralizing each block (three tetrahedra per block). Both outcomes satisfy the same global SALT+MIE constraints, but they differ in $(T,N_{i},S_{i})$, indicating that the multiplicity of internal states at large $S$ is a genuine structural feature rather than noise. Whether this tail multiplicity grows with $V$ (thickening tails) or collapses as $S\to S_{\mathrm{ub}}(V)$ is left to future enumeration and experimentation.
Numerical values of $\mathrm{Combinations}_{\mathrm{lb}}(V)$ for $4\leq V\leq 20$, together with a graphical illustration of its quadratic growth, are given in Appendix E. Notably, this lower-bound count grows only polynomially in $V$, whereas the number of admissible face-type configurations grows combinatorially, as it is controlled by the cumulative sum of integer partition counts $\sum_{S=0}^{S_{\mathrm{ub}}}p(S)$, as elaborated in the following section.
@@
When attempting to uncover the actual face-type and vertex-type configurations corresponding to a specific polyhedron defined by vertex count $V$, triangulation segment count $S$, and tetrahedron count $T$, the procedures described in Section 7 should be applied. These methods allow one to solve for specific polygonal face arrangements and converge to vertex degree structures that are compatible with a given global combinatorial configuration.
@@ 8.2 Estimating the Upper Bound from Flatness and Face Combinations
While previous sections emphasize compatibility between internal and external combinatorics, it is also useful to estimate an upper bound of possible polyhedral configurations based solely on the total surface angle budget. In contrast to a full combinatorial analysis of vertex and edge constraints, this method estimates the number of ways one can arrange polygonal faces on a polyhedral surface using only the flatness range of the surface. It does not take into account local realizability constraints, such as vertex incidence or internal fillability by tetrahedra. As it doesn’t account for vertex types, it serves as a partial initial upper-bound estimate of surface complexity.
@@ 8.3 Method Outline
Define the triangulation segment budget $S$, where $S=0$ corresponds to a surface composed entirely of triangles (i.e., zero extra flatness beyond minimal curvature), and larger values of $S$ correspond to the inclusion of increasingly flat faces (e.g., quadrilaterals, pentagons). The minimum is always $S_{\mathrm{lb}}=0$, corresponding to the fully triangulated case. The maximum possible number of triangulation-equivalent segments $S_{\mathrm{ub}}$ depends on the number of vertices $V$ and follows the previously mentioned parity–sharp linear functions:
$S_{\mathrm{ub}}=\left\{\begin{array}[]{ll}\frac{3}{2}V-\tfrac{13}{2}&\text{if }V\text{is odd}\\ \frac{3}{2}V-6&\text{if }V\text{is even}\end{array}\right$.
@@
Thus, every value of $S\in[0,S_{\mathrm{ub}}]$ corresponds to a number of integer partitions under these constraints. The total number of valid polygonal configurations at a given $V$ is then equal to the sum of these partition counts from $S_{\mathrm{lb}}$ to $S_{\mathrm{ub}}$. If no restriction is placed on the number of polygon sides, then this count becomes equivalent to the unrestricted partition function $p(S)$, a classical object in number theory that enumerates the number of ways to write $S$ as a sum of positive integers (ignoring order) [andrews1976]. The partition function begins:
@@
$\sum_{S=0}^{S_{\mathrm{ub}}}p(S)$
For instance, for $V=5$, I compute $S_{\mathrm{ub}}=1$, and the number of valid configurations is:
@@
Table 3: Lower and upper bounds on total external flatness $S$ and the resulting upper bound on admissible face-type combinations, as a function of vertex count $V$, under the symbolic enclosure constraints.
$V$ $S_{\mathrm{lb}}$ $S_{\mathrm{ub}}$ Face-type combinations (upper bound)
@@
If a maximum polygon side count $k_{\max}$ is imposed (such as limiting the use of hexagons or excluding higher-order polygons) the enumeration becomes a case of restricted integer partitions, where only parts of size $\leq k_{\max}-3$ are allowed. This constraint reflects that each non-triangular face contributes a specific number of excess triangulation segments. Restricted partitions offer a more realistic estimate by excluding configurations that rely on geometrically unlikely face types. For example, limiting the count of hexagons in some configurations prevents overcounting face combinations that cannot physically tile a polyhedral surface. By summing the number of valid restricted partitions for each value of $S=0$ to $S_{\mathrm{ub}}$, I obtain a tighter upper bound on the number of feasible face-type combinations, filtering out many false positives that are combinatorially valid but geometrically unconstructible.